OpenStudy (anonymous):

Solve the equation. 4/3 + 2/3q = -5/3 -q -1/3 Show steps, please! :)

OpenStudy (anonymous):

@dpaInc plzz help her

hero (hero):

\[\frac{4}{3} + \frac{2q}{3} = \frac{5q}{3} - \frac{1}{3}\]

hero (hero):

\[\frac{4}{3} + \frac{1}{3} = \frac{5q}{3} - \frac{2q}{3}\]

hero (hero):

\[\frac{5}{3} = \frac{3q}{3}\]

hero (hero):

\[\frac{5}{3} = q\]

OpenStudy (anonymous):

Thanks so much :)

hero (hero):

yw